My question is:
A rectangle has sides of 3x-4 and 7x+10. Find the expression that represents its perimeter.
Could someone help me with this one and show me how you set it up to find the answer? Thank You.

Answer by checkley75(3666)

The poerimeter of a rectangle is 2L+2W.
Thus 2(3x-4)+2(7x+10)=perimeter
6x-8+14x+20=perimeter
20x+12=perimeter
-----------------------------
20X=PERIMETER-12
X=(PERIMETER-12)/20
-----------------------------
4(5x+3)=perimeter
-----------------------------
THE ONLY MISSING ELEMENT IS EITHER THE PERIMETER VALUE OF THE X VALUE.